A 500 kg cart is rolling to the right at 3.50 m/s. A 70.0 kg man is standing on the right end of the cart.

What is the speed of the cart if the man suddenly starts running to the left with a speed of 10.0 m/s relative to the cart?

Express your answer with the appropriate units.
Answer #1

using conservation of momentum

(m1 + m2) v = m1 v1 + m2 v2

now velocity of cart with respect to ground

v1 = v2 - 10

then

(500 + 70)* 3.5 = 70* (v2 - 10 )+ 500 v2

v2 = 4.728 m/s

=====

Comment before rate in case any doubt, will reply for sure.. goodluck